OpenStudy (nicoleg7):

Will fan and medal Robin can mow a lawn in 3 hours, while Brady can mow the same lawn in 4 hours. How many hours would it take for them to mow the lawn together? seven twelfths twelve sevenths 7 12

563blackghost (563blackghost):

To solve we use... \(\huge\bf{\frac{1}{\frac{1}{3} + \frac {1}{4}}}\)

OpenStudy (nicoleg7):

thxs again @563blackghost